2.A car travels around a bend of radius 400m on a road which is banked at an angle to the horizontal, if the car has no tendency to slip when travelling at 35m/s. Find the angle
1
Expert's answer
2019-11-04T10:12:42-0500
"\tan{\theta}=\frac{\frac{mv^2}{r}}{mg}=\frac{v^2}{rg}"

"\tan{\theta}=\frac{35^2}{(400)(9.8)}=0.3125"

"\theta=\arctan{0.3125}=17.35\degree"